perform the indicated operation and,if possible simplify
32/45 / 48/35=
55/86 - 27/86=

5/11 + 1/4=

Answer by stfuenaira(10)

1.
32/45 divided by 48/35
32/45 times 35/48
Answer: 14/27
2.
55/86 - 27-86 = 28/86 = 14/43
Final Answer: 14/43
3.
5/11 + 1/4
= 20/44 + 11/44
Final Answer: 31/44